OpenStudy (anonymous):

write an equation in the form y=mx+b for the line that passes through the point (-5,-1) and has a slope of -2

OpenStudy (anonymous):

So what do you think m would be?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

so write y=-2x + b, then sub in x= -5 and y=-1

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

-1 = -2(-5) + b, so -1=10+b so -11=b :)

OpenStudy (anonymous):

so its -1=10=-11 the answer

OpenStudy (anonymous):

yeah, b= -1 - 10 = -11

OpenStudy (anonymous):

im confused is the y and the x show

OpenStudy (anonymous):

x = -5 and y= -1 right?

OpenStudy (anonymous):

from the point (-5, -1)

OpenStudy (anonymous):

\[y=-2x + b\] \[-1 = -2(-5) + b\] \[-1=10 + b\] \[-11 = b\]